• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 10:52:30 on April27,2008

    Entry number 235092

    keyword=Setting up for kinematics L

    10:40

    Todd moved to LHRS=12.5
    Brad tied up a cable to red light fixture so it would not get hung up in the HRS movement

    Changed ethan/argon tank. Set to 12 psi
    Did not change Helium tank.
    Balloons have not been filled in 64 hours. Too Long.
    BigBite spec balloon sagging badly away from pole faces. at least 10 cm on each side at about 20 cm into pole face.
    MWDC balloon also sagging.
    Filled both balloons
    Return to counting house

    Setting up for kinematics L
    Recylced Q2 and Q3 to 1400 A
    Set P=1.194
    Beam cuurent = 3 ua
    Set prescales ps1=3600,ps2=11800,ps3=4,ps4=1,ps5=1,ps6=1,ps7=330,ps8=100

    Request MCC to Mask target
    Run 4110 Tantalum elastic done 0.5M events
    Run 20781 8500 events

    Request MMC to Mask target
    Change to carbon target